bella2025:

4. solve by Substitution -3x + 2y = 11 x = y + 4

bella2025:

-3x+2(1)=11 am I right

bella2025:

-3x+2=11 -2 -3x=9/-3 x=-3

Joe348:

so what you want to do is plug in the second equation into the first equation

Joe348:

it would look like this

bella2025:

@joe348 wrote:
it would look like this
I see what I did wrong I was trying to solve for y first instead of x

bella2025:

@joe348 wrote:
it would look like this
-3(1y+4)+2y=11

Joe348:

no

Joe348:

it should be \[-y-12=11\]

Joe348:

then add 12 on both sides

bella2025:

where do you get 12

Joe348:

\[-y=23\]

Joe348:

@bella2025 wrote:
where do you get 12
I substitued the second equation for the first equation

Joe348:

so like this ...

Joe348:

\[-3(y+4)+2y=11\]

Joe348:

then i simplied it with \[-y-12=11\]

Joe348:

then on both sides you add 12

Joe348:

\[-y=23\]

bella2025:

@joe348 wrote:
\[-y=23\]
where did you get -y

Joe348:

then divided by that -y and you get \[-23\]

Joe348:

@bella2025 wrote:
@joe348 wrote:
\[-y=23\]
where did you get -y
okay you see the 2nd equation plugged in?

Joe348:

after that i simplified the equation

Joe348:

like i was saying \[y=-23\]

Joe348:

so since \[y=-23 \] you want to plug that in for the second equation to figure out what X is

Joe348:

so you know the second equation is \[x=y+4\] so you plug \[-23 \] for y

Joe348:

so \[x=-23+4\] then subtract that and you should get \[x=19\]

Joe348:

so your final answer are \[x=19 \ and\ y=-23\]

Joe348:

did i help in any way or are you still confused?

bella2025:

@joe348 wrote:
did i help in any way or are you still confused?
no I figured it out

Joe348:

oh okay good